OpenStudy (anonymous):

Anyone know how to solve for y in this equation (1/2)e^2y=x^3+(3/8)

OpenStudy (tkhunny):

Perhaps introduce a logarithm?

OpenStudy (anonymous):

when u introduce the log u get ln(e^2y)=ln(x^3+3/8)/1/2 then u get 2y= ln(8x^3+3)/4 that does not make sense though i think you check the variable it has to be solved by log but you cant do that if you have the different variables on both sides

OpenStudy (anonymous):

A solution is attached. Mathematica was used for most of the calculations.